OpenStudy (anonymous):

If f(x)=2⋅3^x, which of the following is equal to f(1+x)-f(x)? 2f(x) b. (1+x)f(x) c. -3f(x) d. f(1)

jimthompson5910 (jim_thompson5910):

f(x)=2*3^x f(1+x)=2*3^(1+x) f(1+x)=2*3^1*3^x f(1+x)=2*3*3^x f(1+x)=6*3^x ------------------------------ f(1+x)-f(x) 6*3^x - 2*3^x (6-2)*3^x 4*3^x 2*2*3^x 2*(2*3^x) 2*f(x) So f(1+x)-f(x) = 2*f(x) where f(x) = 2*3^x

OpenStudy (anonymous):

THANK YOU SO MUCH @jim_thompson5910 . got it!

jimthompson5910 (jim_thompson5910):

yw